### v1

# On the Adversarial Robustness of LASSO Based Feature Selection

Abstract
In this paper, we present a novel feature selection method based on the Wasserstein distance. Feature selection plays a critical role in reducing the dimensionality of input data, thereby improving machine learning efficiency and generalization performance. Unlike traditional feature selection approaches that rely on criteria such as correlation or KL divergence, our method leverages the Wasserstein distance to measure the similarity between distributions of selected features and original features. This approach inherently accounts for similarities between classes, making it robust in scenarios involving noisy labels. Experimental results demonstrate that our method outperforms traditional approaches, particularly in challenging settings involving noisy labeled data.
Index Terms:
Feature selection, Wasserstein distance, label correlation, noisy label, graphical model.
 [^0] This work was accomplished at UC, Davis during Fuwei Li’s (e-mail: fuweily@gmail.com) Ph.D. study.
I Introduction
Feature selection is a thoroughly studied topic in recent decades. In the preliminary phase of this research topic, it is aimed to increase the performance of the machine learning algorithm by selecting the most important features. At that time, machine learning algorithms are far less powerful than today’s, so by reducing the dimension of the input data features, the algorithm can learn much faster and more accurate. Also, the limited storage capacity at that time required the data samples to be as small as possible while keeping the most informative features. Now, we have much more powerful and efficient machine learning algorithms. For example, the deep neuron networks can handle billions of data samples. It seems that there is no need to do feature selection anymore. However, the data generated by humans and the Internet of Things(IoTs) have also increased dramatically. The rate grows even faster than the storage capacity as well as the computing power. In the near future when the IoTs play a more important part in our life, the data generated by IoTs and humans will be expected to grow exponentially.
Even though we have larger storage capacity and greater computing power, there are still a lot of potential benefits to doing feature selection. First, feature selection can speed up the computation of machine learning. There are two phases in today’s machine learning algorithm, the training phase and the test phase. In the training phase, the parameters of the learning machine grow with the dimension of the input data. For example, the deep neuron network, facilitated by hundreds of network layers and thousands of layer width, while making the training and inference very slow. If the dimension of the input feature is reduced, it will make learning machines more efficient and speed up the training and inference process. Second, it can help us get better generalization performance through reducing the dimension of the input data. It is proved that with fixed number of samples, the more powerful of the learning machine, the more likely it will overfit the data and thus poorer generalization performance. Even the data dimension is high, only small parts of it contains the most useful information [1]. By feature selection, we can distinguish those important features among the high dimensional features of the data, use simpler learning machines, and get better generalizations. Third, it can also reduce the storage capacity to low the cost. Though the price of the storage device reduce almost linearly over years[2], the data grows exponentially. If the dimension of the data is reduced, it will be economically efficient.
Algorithms of feature selection can be divided into three categories[3]: wrapper, embed, and filter methods. Wrapper methods use some exhaustive searching among different subsets of features and evaluate the feature sets based on a specific learning machine. These methods lead to very high computational complexity and are feasible only when the number of the features is small, it is seldom used nowadays.
The embed methods integrate feature selection into the design of the learning machine. Therefore, feature selection is accomplished since the training of the learning machine. The embed method includes various kinds of spars driven algorithms[4][5]. Since feature selection is done together with the training, they are time-saving. However, these methods require careful designing. Different learning algorithms and different parameters may lead to totally different features.
The filter method is trying to design a scheme independent of the learning machines. Hence, compared with the wrapper and embed methods, it is suitable for broader kinds of learning machines. So, in this paper, we will focus on the filter method.
Firstly, it is necessary to clarify the feature selection problem. Suppose there are a bunch of data samples $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$, where $N$ is the number of samples, $\mathbf{x}_{i}$ is the feature vector for sample $i$, and $y_{i}$ is the label for sample $i$. $\mathbf{x}_{i}\in\mathbb{R}^{M}$ is a $M$ dimensional vector and $\mathbf{x}_{i}=[x_{1}^{i},x_{2}^{i},\cdots,x_{M}^{i}]$. Each component of $\mathbf{x}_{i}$, $x_{j}^{i}$, is called a feature of samples $i$. For single labeled data, the label $y_{i}$ belongs to one of the class labels $y_{i}\in\{c_{1},c_{2},\cdots,c_{n_{c}}\}$ and for multi labeled data, the label $y_{i}$ belongs to subset of the class labels $y_{i}\subseteq\{c_{1},c_{2},\cdots,c_{n_{c}}\}$, where $n_{c}$ denotes the number of classes. Further, all the samples are assumed to be independent and identically distributed according to distribution $p(\mathbf{X},Y)$. We also use the set $\theta\subseteq\{1,2,\cdots,M\}$ to represent the index of the feature we are choosing and $\mathbf{X}_{\theta}$ denotes the selected features which are indexed by $\theta$.
Various kinds of criteria are proposed to select the best feature set, including the correlation based criterion[6], probability measure based criterion[7], and conditional likelihood based criterion[8]. Since most of the machine learning algorithms are intend to use a parametric model to mimic the empirical data distribution under some probabilistic distance measure, it is natural to choose the probability measure as the feature selection criterion. For a learning problem, given an instance of the feature $\mathbf{X}=\mathbf{x}$, we want to know the distribution of the classes. This class conditional distribution can further used to construct an optimal Bayesian estimator. So, the class conditional distribution is the most important part in a learning problem. After feature selection, we are only given a subset of the feature, $\mathbf{X}_{\theta}=\mathbf{x}_{\theta}$, we still intend to make the class distribution be similar to the original class distribution given the whole features. So, we minimize the distance between the two conditional distributions, the class conditional distribution given all the features $p(Y|\mathbf{X}=\mathbf{x})$ and the class conditional distribution given the selected features $p(Y|\mathbf{X}_{\theta}=\mathbf{x}_{\theta})$. Formally, we attempt to solve this minimization problem
$\underset{|\theta|=K}{\text{argmin}}\,D\,[\,p(Y|\mathbf{x}),\,p(Y|\mathbf{x}_{\theta})\,]$
where $D\,[\cdot,\cdot]$ is some probabilistic distance measure to tell the distance between two probability distributions, $|\theta|$ means the cardinality of the set $\theta$ and $K$ is the number of features we want to choose. Given the weights of different instances of feature $\mathbf{X}=\mathbf{x}$, it is better to use the expected value of $D$ as the distance measure. So, our minimization problem is
$\underset{|\theta|=K}{\text{argmin}}\,\mathbb{E}_{\mathbf{X}}D\,[\,p(Y|\mathbf{X}),\,p(Y|\mathbf{X}_{\theta})\,]$ (1)
Later we will detail the choosing of our distance measure $D$ as the Wasserstein distance and explain why we choose that distance measure.
Traditionally, people may use measures induced by information theory, for example, the Kullback-Leibler (KL) divergence, f-divergence, and $\chi^{2}$ divergence. Among the three, the KL divergence is widely used[7][8][9]. Even though KL divergence is simple, it still has some weaknesses. The definition of KL divergence between $p(Y|\mathbf{x})$ and $p(Y|\mathbf{x}_{\theta})$ is $\sum_{i=1}^{n_{c}}p(y=i|\mathbf{x})\ln\frac{p(y=i|\mathbf{x})}{p(y=i|\mathbf{x}_{\theta})}$. It can also be written as an expectation form, that is, $\mathbb{E}_{p(Y|\mathbf{x})}\ln\frac{p(y|\mathbf{x})}{p(y|\mathbf{x}_{\theta})}$. The definition indicates the KL divergence treats all the classes equally. For example, if $p(Y|\mathbf{x})=[0.5,0.5]$ and $p(Y|\mathbf{x}_{\theta_{1}})=[0.1,0.9]$, and another distribution is $p(Y|\mathbf{x}_{\theta_{2}})=[0.9,0.1]$, then we have $D[p(Y|\mathbf{x}),p(Y|\mathbf{x}_{\theta_{1}})]=D[p(Y|\mathbf{x}),p(Y|\mathbf{x}_{\theta_{2}})]$. Even if we have two totally different distributions $p(Y|\mathbf{x}_{\theta_{1}})$ and $p(Y|\mathbf{x}_{\theta_{2}})$, the KL distance measure can not distinguish them.
Figure 1: The first column indicates the original conditional distribution for different classes. The second column indicates the conditional distribution based on feature set $\theta_{1}$. The third column shows the conditional distribution based on feature subset $\theta_{2}$. If we use KL distance as the features selection criterion, the two selected feature subset will lead to the same distance. So, there will be no differences if we select those two feature sets. However, intuitively, the second feature subset will give us more reasonable results because the husky is more similar to the alaska than the cat.
In the computation of KL distance, the denominator in the log term also makes the KL divergence unstable. Consequently, when the denominator is approximate to zero, the KL divergence will explode. This property makes feature selection based on the KL distance not stable especially when some noise is introduced into the data samples.
One reason why KL distance is so widely used is its simple and decomposable computation. With this decomposable computing advantage, we can design efficient machine learning algorithms. However, it fails to capture the relationship between different classes. Relationships between classes are very common in real practice. For example, in the text classification problem, one article is assigned to several classes which have hierarchical structures[10]. In ImageNet[11], some of the image labels have very high correlation. Several papers have used this correlation in classification problems[12][13][14]. Nevertheless, little work has done in the context of feature selection. In this work, we assume there is a ground metric, a prior, describing the distances between classes. Here is a simple example to show how useful of this inter-class information. Suppose there are three classes: cat, husky and alaska. If the distribution based on the full features is $p(Y|\mathbf{x})=[0,1,0]$ and two distributions based on different sub features $p(Y|\mathbf{x}_{\theta_{1}})=[1,0,0]$ and $p(Y|\mathbf{x}_{\theta_{2}})=[0,0,1]$. In this case, we would better choose $\mathbf{x}_{\theta_{2}}$ as the selected feature due to the fact that misclassifying a dog as a cat is more significant than as another kind of dog.
Wasserstein distance is also called Earth Movement distance[15]. It is a probability distance measure which inherently uses the class distance. Wasserstein distance is used in many applications including computation graphics[16], clustering[17], Generative Adversarial Networks[18], and so forth. In the latter section, we will detail the computation of Wasserstein distance and how it is used in our feature selection problem.
Another common problem many machine learning tasks encounter is the label noise. It is known that the label noise will dramatically decrease the classifier performance[19]. Sources of label noise may come from various ways. First, label noise may simply come from the transmission process. For example, if the class labeling work is distributed to the workers by the wireless communication channel, the communication noise may add to the labels[19]. Second, label noise may come from the lack of information. For instance, to tell whether a person is male or female, providing his/her voice data is not sufficient. In medical diagnosis, a doctor usually uses several reports provided by different inspection methods to determine a person has a certain kind of illness. Third, labels may come from some unreliable sources. In particular, now we have billions of data need to be labeled and the price to do label by human experts is very high. Instead, people are trying to use some cheap but not reliable source, like Amazon Mechanical Turk[20][21]. Fourth, when the class is subjective, for example, in medical diagnosis, for a certain medical image, different experts may give different results. This happens especially when two classes have some conceptual confusions. So, label noise commonly exists in real practical dataset.
Various methods have been introduced to tackle the label noise problem[22]. For example, the filter based methods firstly detect the wrongly labeled data, correct them, and then use a traditional algorithm to do classification. Others use a specific label noise model to capture how label noise is introduced[23] [24] [25]. Even though so many label noise models are proposed, different models can only capture one certain aspect of the label noise. There is no unified way to model the label noise. In real practice, it is even more complicated, because in most cases, there is no ground truth label noise model. Though there is no unified way to describe the label noise, all of them admit that label noise is more likely to happen between two similar objects. We will show that our Wasserstein distance based algorithm captures this property naturally. Thus our algorithm is more stable and more robust to the label noise.
In the following section, we will first introduce the Wasserstein distance. Next, we will introduce the probabilistic graphic model and show how Markov Blanket has been used in the feature selection problem. Then, we will introduce our feature selection algorithm. Last, we will show the performance of our algorithm in the noisy label settings.
II The Wasserstein Distance
II-A Definition of Wasserstein Distance
Wasserstein distance is also called earth mover’s distance[26] or optimal earth transport distance[15]. Suppose there are two measures $\nu_{1}$ and $\nu_{2}$ on $\mathcal{C}$, and a cost function $d:\mathcal{C}\times\mathcal{C}\to\mathbb{R}$ measures the cost of transporting the measure of a mass on $\mathcal{C}$ to another $\mathcal{C}$. The Wasserstein distance is defined as the cheapest way to transport the mass with the probability measure $\nu_{1}$ to the corresponding mass with the probability measure $\nu_{2}$:
$D_{wass}(\nu_{1},\nu_{2})=\inf_{\pi\in\Pi(\nu_{1},\nu_{2})}\int_{\mathcal{C}_{1}\times\mathcal{C}_{2}}d(c_{1},c_{2})\pi(\mathrm{d}c_{1},\mathrm{d}c_{2})$ (2)
where $\Pi(\nu_{1},\nu_{2})$ is the set of joint probability measure with marginal probability measure $\nu_{1}$ and $\nu_{2}$ respectively. Solving this optimization problem means to find the joint probability which minimizes the cost of transporting one mass with measure $\nu_{1}$ to the corresponding mass with measure $\nu_{2}$. In this paper, we only focus on the discrete measure on the classes. In this case, the Wasserstein distance between probability measure $\mathbf{p}(Y)=[p(y_{1}),p(y_{2}),\cdots,p(y_{n_{c}})]$ and $\mathbf{q}(Y)=[q(y_{1}),q(y_{2}),\cdots,q(y_{c})]$ can be computed by solving the optimization problem:
$\min\quad$ $\text{tr}(\mathbf{Q}^{\top}\mathbf{D})$
s.t. $\mathbf{Q}\cdot\mathbf{1}=\mathbf{p}(Y)$
$\mathbf{Q}^{\top}\mathbf{1}=\mathbf{q}(Y)$
$Q_{i,j}\geq 0,\,\forall\,i,j\in\{1,2,\cdots,n_{c}\}$ (3)
where $\mathbf{Q}$ is the joint probability matrix with marginal distribution $\mathbf{p}(Y)=[p(Y=c_{1}),p(Y=c_{2}),\cdots,p(Y=c_{n})]$ and $\mathbf{q}(Y)=[q(Y=c_{1}),q(Y=c_{2}),\cdots,q(Y=c_{n})]$; $\mathbf{D}$ is the distance matrix with its elements defined as $D_{ij}=d(c_{i},c_{j})$ and $d(c_{i},c_{j})$ defines the distance from class $c_{i}$ to class $c_{j}$. The distance $d(c_{i},c_{j})$ measure the dissimilarity between class $c_{i}$ and class $c_{j}$. This is a linear programming problem. To solve this linear programming, the computational complexity would be $\mathcal{O}(n_{c}^{3.5})$[27]. The computation is heavy when the number of classes is large. Recently, [28] proposed a regularized form of this problem. They added a negative entropy of the joint distribution matrix regularization term in the objective function. Accordingly, we are solving the minimization problem with an objective function as
$\min:\text{tr}(\mathbf{Q}^{\top}\mathbf{D})-\frac{1}{\lambda}h(\mathbf{Q})$
where $h(\mathbf{Q})$ computes the entropy of the joint matrix $\mathbf{Q}$ and is defined as $h(\mathbf{Q})=-\sum_{i,j}Q_{ij}\ln Q{ij}$. In this way, the new problem can be solved efficiently by a matrix balance iteration process. This entropy regularized algorithm results in a computational complexity of $\mathcal{O}(n_{c}^{2})$ and maintains high accuracy. In our algorithm, we will use this method to compute the Wasserstein distance.
III Markov Blanket and Feature Selection
Markov Blanket is widely used in feature selection[29][30][31][32]. Markov Blanket is the smallest feature set which contains all the information about the target. In a directed probabilistic graphic model, the Markov Blanket of a node is the nodes’ parents, children and spouses. Figure(2) shows a probabilistic graphic model, where the Markov Blanket of node A contains the nodes in the shaded area. More formally, the Markov Blanket of the target $Y$ is the smallest set $\mathcal{B}\subset\{1,2,\cdots,M\}$ which have $p(Y=y|\mathbf{X}=\mathbf{x})=p(Y=y|\mathbf{X}_{\mathcal{B}}=\mathbf{x}_{\mathcal{B}})$ for every instance of $\mathbf{X}$, where the elements of $\mathcal{B}$ indicate which features are chosen.
Figure 2: The probabilistic graphic model, the Markov Blanket of node A contains the nodes in the shaded area.
Ideally, if we want to do feature selection without information loss, the Markov Blanket of the target is the best feature set we can get. Most of the features selection algorithms based on Markov Blanket mentioned above utilize the KL divergence distance. In this paper, we are going to introduce the Wasserstein distance into the feature selection. So, we will start from the basic feature selection algorithm described in [29]. Then, we will show the benefits of using the Wasserstein distance in feature selection.
To further explain our algorithm, we need elaborate some independence properties under the Wasserstein distance measure.
Proposition 1 $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$ iif $D_{wass}[p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}),p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})]=0$. where $D_{wass}[\mathbf{p},\mathbf{q}]$ denotes the Wasserstein distance between distribution $\mathbf{p}$ and $\mathbf{q}$.
Proof: Suppose the elements of the distance matrix $\mathbf{D}$ in Problem(3) are positive except the diagonal elements and let $\mathbf{p}(Y)=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}),\mathbf{q}(Y)=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. The solution of the optimization problem(3) is zero if and only if the joint distribution matrix $Q_{i,j}=0$ for all $i\neq j$. Then we have $\mathbf{Q}\cdot\mathbf{1}=\mathbf{Q}^{\top}\mathbf{1}$, so $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. Also, if $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. It is easy to see that the optimal solution of problem(3) is zero.
Proposition 1 tells us that the Wasserstein distance can reveal the independence property in the probabilistic graphic model.
In the paper [7], they tried to find a Markov Blanket for each feature $X_{i}$. However, finding the Markov Blanket for a feature is too hard in practice. Proposition 2 tells us a more practical way to find whether $X_{i}$ belongs to the Markov Blanket.
Proposition 2 Let $\mathbf{X}_{\mathcal{G}_{i}}$ denotes a subset of $\mathbf{X}\setminus X_{i}$ and contains all the features which $X_{i}$ directed to. If for every setting of $Y=y,\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}$ we have $p(Y=y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y=y|\mathbf{X}_{\mathcal{G}_{i}})$. Then, $X_{i}$ does not belong to the Markov Blanket.
Proof: Since $X_{i}$ and $Y$ are independent given $\mathbf{X_{\mathcal{G}_{i}}}$ and $\mathbf{X}_{\mathcal{G}_{i}}$ contains all the nodes which $X_{i}$ direct to, if $X_{i}$ belongs to the Markov Blanket, there should be a node which both $Y$ and $X_{i}$ directed to, which contradicts the assumption that $Y$ and $X_{i}$ are independent given $\mathbf{X}_{\mathcal{G}_{i}}$.
As [7] has pointed out, even if we can find the appropriate feature set $\mathbf{X}_{\mathcal{G}_{i}}$, in practice, the distance of the two conditional probabilities,$p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})$ and $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$ , will hardly be zero. This is due to the noise or the limited samples. To test whether $p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}})$, we need compare every configuration of $\mathbf{X}_{\mathcal{G}_{i}},X_{i}$. To make it more convenient and take the weights of the $\mathbf{X}_{\mathcal{G}_{i}},X_{i}$ into consideration, we use the averaged distance to measure their discrepancy. The averaged distance is defined as:
$\delta(X_{i})=\mathbb{E}_{\mathbf{X}_{\mathcal{G}_{i}},X_{i}}D_{wass}[p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i}),p(Y|\mathbf{X}_{\mathcal{G}_{i}})]$ (4)
So, if there is a $\mathcal{G}_{i}$ which has $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$ for every $\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}$, $\delta(X_{i})$ should be zero or at least small in practice.
To find an appropriate $\mathcal{G}_{i}$ for feature $X_{i}$, we also use the approximation method used in [7]. We find the top-L most relevant features with $X_{i}$ as $\mathbf{X}_{\mathcal{G}_{i}}$. So, by choosing the most relevant $L$ features as $\mathbf{X}_{\mathcal{G}_{i}}$, probably $\mathbf{X}_{\mathcal{G}_{i}}$ will include all the nodes inbound to or outbound from $X_{i}$ From the probabilistic graphic model, those nodes will cut all paths from $X_{i}$ to $Y$. According to proposition 2 and the definition of $\delta(X_{i})$ in Eq.(4), if $\delta(X_{i})=0$, $X_{i}$ does not belong to the Markov Blanket, we can safely delete feature $X_{i}$.
Here is our algorithm:
  1. 1.
    Compute the absolute value of correlation $\rho_{i,j}=|\frac{Cov(X_{i},X_{j})}{D(X_{i})D(X_{j})}|$ for each pair of feature $X_{i}$ and $X_{j}$. Let $\mathbf{\bar{X}}=\mathbf{X}$.
  2. 2.
    For every $X_{i}\in\mathbf{\bar{X}}$, find the features with $L$ largest value of $\rho_{i,j}$ and define them as $\mathbf{X}_{\mathcal{G}_{i}}$.
  3. 3.
    Compute $\delta(X_{i})$ for each $X_{i}\in\mathbf{\bar{X}}$.
  4. 4.
    Find $X_{i}$ which has the smallest $\delta(X_{i})$, delete $X_{i}$ from $\mathbf{\bar{X}}$, and let $\mathbf{\bar{X}}=\mathbf{\bar{X}}\setminus X_{i}$.
  5. 5.
    Return to (1) until some stopping criteria are satisfied.
IV Feature Selection with Noisy Labels
In this section, we will analyze the performance of our feature selection algorithm based on Wasserstein distance under the noisy labeled data samples. In real practice, the observed samples always contain noisy labels and we can only do feature selection based on these noisy labeled data samples. However, we want to know how good the selected features are when compared with the features selected using the clean data samples.
Label noise exists in most of the data set. Especially, label noise happens between similar classes. Suppose the original data set samples are $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$ and they are drawn independent and identically from the joint distribution $p(\mathbf{X},Y)$. However, because of the label noise what we observed is the noisy labeled data samples, $\{\mathbf{x}_{i},\tilde{y}_{i}\}_{i=1}^{N}$, and we assume they are drawn independent and identically from the joint distribution $p(\mathbf{X},\tilde{Y})$. Here, we assume only the labels of the data samples differ from the original data samples. So, the marginal distributions of features are the same. Since we can only access to the noisy labeled data samples, to select the optimal features, instead of solving the optimization problem Prob.(1), we can only select the optimal features by solving this optimization problem
$\underset{|\theta|=K}{\text{argmin}}\,\mathbb{E}_{\mathbf{X}}D_{wass}[p(\tilde{Y}|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X}_{\theta})]$ (5)
where $D_{wass}[\mathbf{p},\mathbf{q}]$ is Wasserstein distance between $\mathbf{p}$ and $\mathbf{q}$. Here, unlike most of the analysis, we do not assume a specific label noisy model[33] in our analysis. We still use the set $\theta\subseteq\{1,2,\cdots,M\}$ to denote the selected feature set. Fig.(3) shows the relationships between the conditional distributions before and after feature selection in the original clean data set as well as in the observed noisy labeled data set.
Figure 3: The relationships between conditional distributions before feature selection and after feature selection in the original and observed data set respectively
Before further analysis, we need define some quantities to simplify the notation of our analysis. First, $\epsilon_{1}$ refers to the expected Wasserstein distance between the original conditional distribution and the noisy labeled conditional distribution.
$\epsilon_{1}=\mathbb{E}_{\mathbf{X}}D_{wass}[p(Y|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X})]$
Second, $d_{1}(\theta)$ refers to the expected Wasserstein distance between the conditional distribution before feature selection and after feature selection in the original data samples.
$d_{1}(\theta)=\mathbb{E}_{\mathbf{X}}D_{wass}[p(Y|\mathbf{X}),\,p(Y|\mathbf{X}_{\theta})]$
It tells us if we select the feature set $\theta$ and how far we deviate from the original conditional distribution. Similarly, $d_{2}(\theta)$ means the distance in the noisy labeled data set before feature selection and after feature selection:
$d_{2}(\theta)=\mathbb{E}_{\mathbf{X}}D_{wass}[p(\tilde{Y}|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X}_{\theta})]$
In addition, $\epsilon_{2}$ is defined as the expected Wasserstein distance between the conditional distribution in the original clean data sample and the conditional distribution in the noisy labeled data set after feature selection:
$\epsilon_{2}=\mathbb{E}_{\mathbf{X}_{\theta}}D_{wass}[p(Y|\mathbf{X}_{\theta}),\,p(\tilde{Y}|\mathbf{X}_{\theta})]$
Here, the assumption that the same feature set has been chosen in the original data and the noisy labeled data is a bridge to our further analysis.
Using the above notation, we can rewrite our feature selection optimization Prob.(1) and Prob.(5) simply as:
$\theta_{1}^{*}=\underset{|\theta|=K}{\text{argmin}}\quad d_{1}(\theta)$ (6)
$\theta_{2}^{*}=\underset{|\theta|=K}{\text{argmin}}\quad d_{2}(\theta)$ (7)
$\theta_{1}^{*}$ indicates the optimally selected feature set using the original data and $\theta_{2}^{*}$ indicates the optimally selected feature set using the noisy labeled data. Although $\theta_{2}^{*}$ is the best feature set we can get based on the noisy labeled data set, to give a quantitative evaluation of the optimally selected features based on the noisy labeled data set, we use the quantity
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})$ (8)
as the criterion. Here, $d_{1}(\theta_{1}^{*})$ is the baseline and computed based on the clean original data. The smaller of this quantity is, the better of the selected features are. In the following, we will show that how this quantity is related to $\epsilon_{1}$.
To show the value of $d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})$, first we demonstrate the relationship between $d_{1}(\theta)$ and $d_{2}(\theta)$. Then, we show the relationship between the corresponding optimal selected value of $d_{1}(\theta_{1}^{*})$ and $d_{2}(\theta_{2}^{*})$. Finally we will find the value of $d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})$ is bounded by $4\epsilon_{1}$.
Before feature selection, the expected Wasserstein distance between $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ is $\epsilon_{1}$. If the same feature set $\theta$ is selected, will the distance $\epsilon_{2}$ grow very large? The first conclusion demonstrates this will not happen.
Conclusion 1: For the same selected feature set $\theta$ in the original data and in the noisy labeled data, $\epsilon_{2}\leq\epsilon_{1}$.
$\min\quad&\text{tr}(\mathbf{Q}^{\top}\mathbf{D})\\ \text{s.t.}\quad&\mathbf{Q}\mathbf{1}=\mathbf{p}_{Y|\mathbf{x}}\\ &\mathbf{Q}^{\top}\mathbf{1}=\mathbf{p}_{\tilde{Y}|\mathbf{x}}\\ &Q_{i,j}\geq 0,\forall\,i,j\end{aligned}$ $\min\quad&\text{tr}(\mathbf{Q}^{\top}\mathbf{D})\\ \text{s.t.}\quad&\mathbf{Q}\mathbf{1}=\mathbf{p}_{Y|\mathbf{x}_{\theta}}\\ &\mathbf{Q}^{\top}\mathbf{1}=\mathbf{p}_{\tilde{Y}|\mathbf{x}_{\theta}}\\ &Q_{i,j}\geq 0,\forall\,i,j\end{aligned}$
TABLE I: The left optimization problem calculates the distance from the original data to the noisy labeled data; the right optimization problem calculates the distance from the original data to the noisy labeled data after feature selection.
Proof: To compute $\epsilon_{1}$ which involves solving the left optimization problem of Table(LABEL:table1-two-opts) for each configuration of the random vector $\mathbf{X}=\mathbf{x}$ and to compute $\epsilon_{2}$ involves solving the right optimization problem of Table(I). We denote $\mathbf{p}_{Y|\mathbf{x}}=[p(Y=c_{1}|\mathbf{X}=\mathbf{x}),p(Y=c_{2}|\mathbf{X}=\mathbf{x}),\cdots,p(Y=c_{n}|\mathbf{X}=\mathbf{x})]^{\top}$ and the similar vector notation for $\mathbf{p}_{Y|\mathbf{x}_{\theta}},\mathbf{p}_{\tilde{Y}|\mathbf{x}}$, and $\mathbf{p}_{\tilde{Y}|\mathbf{x}_{\theta}}$.
Suppose we have gotten the optimal solution for the left problem as $\mathbf{Q}(\mathbf{x})$, where the optimal solution is denoted as a function of $\mathbf{x}$. So, the optimal value of the left problem is $\sum_{i,j}Q(\mathbf{x})_{i,j}D_{i,j}$. By taking expectation over the random variable $\mathbf{X}$, we get the expression for the value of $\epsilon_{1}$ to be $\epsilon_{1}=\int p(\mathbf{x})\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j}\,\mathrm{d}\mathbf{x}$.
Another matrix $\mathbf{\tilde{Q}}(\mathbf{x}_{\theta})$ with entries defined as
$\tilde{Q}_{i,j}(\mathbf{x}_{\theta})=\frac{\int p(\mathbf{x})Q_{i,j}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
where $\bar{\theta}$ is the complement of set $\theta$, i.e. $\bar{\theta}=\{1,2,\cdots,M\}\setminus\theta$. In the following, we will show that this matrix satisfies the constraints in the right optimization problem.
First, we use Bayesian rule to calculate $p(Y=c_{i}|\mathbf{x}_{\theta})$:
$p(Y=c_{i}|\mathbf{x}_{\theta})=\frac{\int p(Y=c_{i}|\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$ (9)
According to the first constraint in the left problem, we have:
$p(Y=c_{i}|\mathbf{x})=\sum_{j}Q_{i,j}(\mathbf{x})$
Plug this equation into Eq.(9) and we get
$p(Y=c_{i}|\mathbf{x}_{\theta})$ $=\frac{\int\sum_{j}Q_{i,j}(\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\frac{\int Q_{i,j}(\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\tilde{Q}_{i,j}(\mathbf{x}_{\theta})$
It is clear that $\mathbf{\tilde{Q}}(\mathbf{x}_{\theta})$ satisfies the first constraint of the right optimization problem.
Accordingly, following the same routine above, we get
$p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})=\frac{\int p(\tilde{Y}=c_{i}|\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
Since $p(\tilde{Y}=c_{i}|\mathbf{x})=\sum_{j}Q_{j,i}(\mathbf{x})$, plugging these into the calculation of $p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})$, we get
$p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})$ $=\frac{\int p(\mathbf{x})\sum_{j}Q_{j,i}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\frac{\int p(\mathbf{x})Q_{j,i}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\tilde{Q}_{j,i}(\mathbf{x})$
which indicates the joint distribution matrix $\mathbf{\tilde{Q}}(\mathbf{x}_{\theta})$ also satisfies the second constraint of the right optimization problem. Thus, $\mathbf{\tilde{Q}}(\mathbf{x}_{\theta})$ is a feasible solution of the right optimization problem. So,
$\epsilon_{2}$ $=\mathbb{E}_{\mathbf{x}_{\theta}}D_{wass}[p(Y|\mathbf{X}_{\theta})\,,p(\tilde{Y}|\mathbf{X}_{\theta})]$
$\leq\int\sum_{i,j}\tilde{Q}_{i,j}(\mathbf{x}_{\theta})D_{i,j}p(\mathbf{x}_{\theta})\,\mathrm{d}\mathbf{x}_{\theta}$
$=\int\sum_{i,j}\frac{\int Q_{i,j}(\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x_{\bar{\theta}}}}{p(\mathbf{x}_{\theta})}p(\mathbf{x}_{\theta})D_{i,j}\,\mathrm{d}\mathbf{x}_{\theta}$
$=\int\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j}p(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\epsilon_{1}$
It demonstrates that $\epsilon_{2}$ is always less than or equal to $\epsilon_{1}$. By constraining selecting the same feature set under the original data and under the noisy labeled data, the value of $\epsilon_{2}$ is constrained to be no larger than $\epsilon_{1}$.
Using the triangular inequality property of Wasserstein distance, we get
$-\epsilon_{1}-\epsilon_{2}\leq d_{1}(\theta_{1}^{*})-d_{2}(\theta_{1}^{*})\leq\epsilon_{1}+\epsilon_{2}$ (10)
$-\epsilon_{1}-\epsilon_{2}\leq d_{1}(\theta_{2}^{*})-d_{2}(\theta_{2}^{*})\leq\epsilon_{1}+\epsilon_{2}$ (11)
where $\theta_{1}^{*}$, $\theta_{2}^{*}$ have been defined in equation(6)(7). Since $\theta_{1}^{*}$ is the optimal solution for $\min_{|\theta|=K}d_{1}(\theta)$ and $\theta_{2}^{*}$ is the solution for $\min_{|\theta|=K}d_{2}(\theta)$, we also have
$d_{1}(\theta_{2}^{*})\geq d_{1}(\theta_{1}^{*})$ (12)
$d_{2}(\theta_{1}^{*})\geq d_{2}(\theta_{2}^{*})$ (13)
Combine inequalities (10)(11)(12)(13) and we have
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})+d_{2}(\theta_{1}^{*})-d_{2}(\theta_{2}^{*})\leq 2\epsilon_{1}+2\epsilon_{2}$ (14)
Since $d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})\geq 0$, $d_{2}(\theta_{1}^{*})-d_{1}(\theta_{2}^{*})\geq 0$ and $\epsilon_{2}\leq\epsilon_{1}$, we have
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})\leq 4\epsilon_{1}$ (15)
This inequality shows that the optimal feature set $\theta_{2}^{*}$ under noisy labeled data is $4\epsilon_{1}$ suboptimal in the original clean data.
However, we should notice that the value $\epsilon_{1}$ is fixed for a given data set and it is determined by the physical system. Although we can not change this value, we wish this value to be small. If this value is large, which means the data set quality is bad, it can not provide us enough information to select useful feature set. If we can get an estimated value of $\epsilon_{1}$ and get further analysis, it would be greatly helpful. In the following, we will give some insights into the value of $\epsilon_{1}$ and show this value will not be too large.
To compute $\epsilon_{1}$, we need to solve the optimization problem in the left of Table(I) for every configuration of $\mathbf{X}=\mathbf{x}$. Because this optimization problem does not have a closed form solution in general cases, we can not get a closed form solution for $\epsilon_{1}$. Also, the relationship between $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ is not specified, which makes this problem even harder to solve. Instead of directly solving this problem, we can find some meaningful bound for $\epsilon_{1}$.
There are two most widely used noisy label models, which are noise at random (NAR) and noise not at random (NNAR)[22]. NAR model makes assumption that the observed label depends on the underlying true label. So, the observed class conditional distribution given the feature $\mathbf{X}$ is:
$p(\tilde{Y}|\mathbf{X})=\sum_{Y}p(\tilde{Y}|Y)p(Y|\mathbf{X})$ (16)
where $p(\tilde{Y}|\mathbf{Y})$ denotes the transition probability which quantifies the the probability that the true label is $Y$ while observed is $\tilde{Y}$. Under this model, it is still hard to get closed form solution of $\epsilon_{1}$. However, we can construct a joint distribution matrix and get it’s upper bound. Let the elements of joint distribution matrix be $Q_{i,j}(\mathbf{x})=p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})$ and $p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})=p(\tilde{Y}=c_{j}|Y=c_{i})p(Y=c_{i}|\mathbf{x})$. We can verify that this matrix is a feasible solution of the left problem of Table(I). So, we have
$\epsilon_{1}$ $\leq\int p(\mathbf{x})\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j},\mathrm{d}\mathbf{x}$
$=\int p(\mathbf{x})\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})D_{i,j}\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$ (17)
where $p(Y=c_{i},\tilde{Y}=c_{j})$ indicates the joint probability that the true label is $c_{i}$ while the observed is $c_{j}$. If two classes, $c_{i},c_{j}$, are similar, we should expected that the transition probability $p(\tilde{Y}=c_{j}|Y=c_{i})$ should be large and so is the joint probability $p(Y=c_{i},\tilde{Y}=c_{j})$.
The NNAR model assumes that the observed label depends on both the true label and its features. So, the observed class conditional distribution given the features is:
$p(\tilde{Y}|\mathbf{X})=\sum_{Y}p(\tilde{Y}|Y,\mathbf{X})p(Y|\mathbf{X})$ (18)
Compared with Eq.(16), it only differs in the form of the transition probability. Construct a feasible joint probability matrix as $Q_{i,j}(\mathbf{x})=p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})$ and $p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})=p(\tilde{Y}=c_{j}|\mathbf{Y}=c_{i},\mathbf{x})p(Y=c_{i}|\mathbf{x})$. Under this noise model, we can also find an upper bound of $\epsilon_{1}$:
$\epsilon_{1}$ $\leq\int p(\mathbf{x})\sum_{i,j}p(Y=c_{i},Y=c_{j}|\mathbf{x}),\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$ (19)
This upper bound has the exact same form as the upper bound induced by NAR model except the way of computing the joint distribution.
Generally, let $p(Y,\tilde{Y}|\mathbf{X})$ denotes the joint conditional distribution of $Y$ and $\tilde{Y}$ given feature $\mathbf{X}$, where $Y$ is the original label and $\tilde{Y}$ is the noisy label and it satisfies the constraints of the optimization problem, that is, its marginal conditional distributions are $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ respectively. Then we get an upper bound for optimal value of the optimization problem:
$\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})D_{i,j}$
By taking the expected value regarding $\mathbf{X}$, we get the upper bounded solution of $\epsilon_{1}$:
$\epsilon_{1}$ $\leq\int p(\mathbf{X}=\mathbf{x})\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})D_{i,j}\,\mathrm{d}\mathbf{x}$
$=\sum_{i,j}D_{i,j}\int p(\mathbf{X}=\mathbf{x})p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$ (20)
Different from other algorithms whose bound analysis are based on specific noise label model, this bound only depends on the distance matrix and the joint distribution of original label and noisy label.
To evaluate the bound, we need look inside the bound, the term $p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$. This term involves the joint probability of the original label and noisy label and the distance matrix. If we admit that the larger distance between two classes is, the lower of their joint probability, that is to say, if the distance is large, the joint probability $p(Y=c_{i},\tilde{Y}=c_{j})$ should be small, and vice versa, the term would not be too large. Other information measure based algorithms, like the KL-distance-based algorithm, will result in exploding value of the distance from original to noisy labeled. For example, when calculate the KL distance of the two conditional probabilities, $\sum_{i}p(Y=c_{i}|\mathbf{x})\frac{p(Y=c_{i}|\mathbf{x})}{p(\tilde{Y}=c_{i}|\mathbf{x})}$ and if the noisy label lead to $p(\tilde{Y}=c_{i}|\mathbf{x})$ approximate to zero, we would get the KL distance exploded. Then the algorithm based on KL distance will not provide useful information for how far the noisy labeled data deviates from the original data.
V The algorithm complexity analysis
Suppose there are $N$ samples and $n_{c}$ classes, the number of features is $M$, the number of conditioned features is $L$ and $K$ is the number of features we want to choose. Below we will analyze the computational complexity of the algorithm in each step.
  1. 1.
    Compute the feature correlation matrix. For computing each pair of the feature correlation, the complexity is $N$. Totally we have $\frac{M(M-1)}{2}$ pairs of features. So, the computational complexity is $\mathcal{O}(NM^{2})$;
  2. 2.
    Sort the correlation matrix. For each feature, we would find the most $L$ significant features. The complexity of partial sorting algorithm to find the most relevant $L$ features is $\mathcal{O}(M+L\log L)$. Then the overall complexity of computing the correlation and sorting is $\mathcal{O}(M^{2}(N+M+L\log L))$;
  3. 3.
    To eliminate each feature, we need to calculate the Wasserstein distance between two conditional distributions. The complexity of estimating the conditional probability of $p(Y|\mathbf{X}_{\mathcal{G}_{i}})$ varies with different estimation algorithms. Here we assume all the features are discrete valued. Suppose the features are binary valued, and then the complexity to compute the conditional distribution is $2^{L}N$. The second step is to compute the distance between those two conditional probabilities. Computing the distance involves solving a linear programming problem. The complexity of the linear programming is $\mathcal{O}(n_{c}^{3})$. By using an approximate method in [28], the complexity can be reduced to $\mathcal{O}(n_{c}^{2})$. So, the complexity of calculating the Wasserstein distance of the conditional distribution is $\mathcal{O}(2^{L}Nn_{c}^{2})$;
  4. 4.
    Suppose, we are in the $i$’th step, then there are $M-i$ features left. Thus, we will need $M-i$ times of computing the Wasserstein distance. Then the complexity of eliminating process is $\mathcal{O}((2M-K)K2^{L}Nn_{c}^{2})$. In each eliminating step, we assume re-computing all the $\delta(X_{i})$ for each remaining feature $X_{i}$. Actually, the eliminated feature in the previous step only effect only a small part of the conditional distributions. Then the first term can be reduced to a constant.
In summary, the computational complexity of the algorithm is
$\mathcal{O}(M^{2}(N+M+L\log L)+(2M-K)K2^{L}Nn_{c}^{2})$
When $K$ is small the complexity is approximately to
$\mathcal{O}(M^{3}+M^{2}N+M^{2}2^{L}Nn_{c}^{2})$
When the number of samples is far larger than the number of features, the complexity of the algorithm will be dominated by the term $NM^{2}2^{L}n_{c}^{2}$.
VI Simulation Results
In the first experiment, we use the Enron email corpus data set[34], and some pre-processing had been done in [35]. The data set contains 1648 samples. Each sample includes a 1001 dimensional feature vector $\mathbf{x}_{i}$ and multiple labels $\mathbf{y}_{i}$. Each dimension of the feature indicates one word in the email. Based on the preprocessing in [35], if a word appears in the email, then its corresponding feature will set to be one. So, each feature is binary valued. The labels of the samples indicate the folders where the email belongs to. For example, the label ”inbox/primary/discussion” means the email in the inbox folder and primary sub-folder and discussion sub-sub-folder. So, the label itself has a hierarchic structure.
Figure 4: The three pictures from the left to the right show the hierarchic structure of the labels, the weight assigned to the path between hierarchic labels, and one example to compute the class distance.
The picture on the left of Figure(4) shows one example of the label structure. In this picture, the root node is the label ”inbox”. However, there is no root node in the Enron dataset. In the experiment, we virtually assign a root node in this hierarchic structure. Then, the label structure can be depicted as a tree. The nodes of the tree indicate the class labels. One important thing to be notified is that not only the leaves of the tree represent labels, but also the non-leaf nodes can represent labels. With this tree structure, it provides us a natural way to compute the distance between different classes. The distance between class $i$ and class $j$ is denoted as $D_{ij}=d(c_{i},c_{j})$. To compute the distance between classes in our case, we should first assign weights to the paths between classes. Here we assume the nearer to the root is, the larger weights should be assigned to the paths. The philosophy behind this is that the higher the layer of the nodes is, the larger the disparity of the classes is. Then we can define the class distance, $d(c_{i},c_{j})$, as the shortest path from class $i$ to class $j$. The second picture shows one possible way to assign the weights to the paths. The third picture in Figure(4) shows how to compute the distance between class $4$ and class $5$. Clearly, under this definition, the distance satisfies the triangular inequality property, i.e. $d(c_{i},c_{j})\leq d(c_{i},c_{k})+d(c_{k},c_{j})$.
The Enron dataset labels have a three-layer structure. From the root layer to the third layer, we assign the weights of the paths to be $0.5,0.2$, and $0.05$ respectively. Even though Enron data is a multi-label dataset, we treat it as single labeled data set by doing the following pre-processing. For each sample with multiple labels in the training dataset, we assign the same features to each individual labels and treat it as an independent sample. Using the feature selection algorithm on the training dataset, we can get our selected feature set index $\theta$. Then based on the selected feature set, we retain the features of the training and testing indicated by the feature set. Using the reduced training data, we train a classifier and use this classifier to predict the labels for the testing data set. Here, in our experiment, we use the k nearest neighbor (kNN) classifier. We use the kNN classifier because the classifier makes few assumptions on the data set and has no hyper-parameters needing to be tuned.
To evaluate the results, we use the top-k loss. It is defined as:
$loss=\frac{1}{k}\sum_{i=1}^{k}\min_{y\in\mathcal{L}}d(\hat{y}_{i},y)$
where $\mathcal{L}$ is the true labels and $\hat{y}_{i}$ is the predicted label with the $i^{th}$ highest prediction probability. In this experiment and the following experiments, we set the number of conditional features $L$ to be $3$.
Figure 5: The top-k loss with different k and different number of features.
Figure(5) shows the top-k loss with different k and different numbers of features. Our algorithm is denoted as ’Wass’ while the KL-distance-based algorithm is denoted as ’KL’. From the figure we can see, our Wasserstein -distance-based algorithm has smaller top-k loss in most cases, especially when k is large. And the top-k loss of the two algorithms both increase as k increase.
Figure 6: The left picture shows top-5 loss with different level of label flipping noises and different numbers of features in two algorithms; the right picture shows the averaged loss over different flipping noises described in the left picture.
In the next experiment, we will show the robustness of our algorithm under noisy labels. In this experiment, we randomly flip the labels in the training data set to one of its nearest neighbors with probability $P$, where the nearest neighbors are defined as the labels which have the class distance less than or equal to 0.2. We use feature selection algorithms on this noisy labeled training data and get the selected feature set. Like in the previous experiment, we reduce the feature dimension for the training and testing dataset. Then we use the feature-reduced training data with the right labels to train a classifier, and then we use the classifier to predict the testing data labels.
Figure(6) demonstrates the top-5 losses under different noise levels and different numbers of features using these two algorithms. The left picture shows our algorithm has smaller top-5 losses and behave much more stably than the KL-distance-based algorithm. Then, to make the left picture more clear, we average the top-5 losses over different $P$. This picture better shows our algorithm has lower top-5 losses. For KL distance algorithm, the behavior is a little pathologic. With the number of features increase, the KL distance based algorithm’s top-5 losses even increase when the number of features is over 120. This is due to the fact that under the noisy setting, the KL-distance-based algorithm can not learn the most useful features.
TABLE II: 20 newsgroups dataset topics
comp.graphics
comp.os.ms-windows.misc
comp.sys.ibm.pc.hardware
comp.sys.mac.hardware
comp.windows.x
rec.autosrec.motorcycles
rec.sport.baseball
rec.sport.hockey
sci.crypt
sci.electronics
sci.med
sci.space
misc.forsale
talk.politics.misc
talk.politics.guns
talk.politics.mideast
talk.religion.misc
alt.atheism
soc.religion.christian
Figure 7: top-5 loss with different number of features on the Flickr data set.
Another dataset we have used is the Yahoo/Flickr Creative Commons 100M dataset[36][12]. It is a multi-label dataset. Each sample is an image with some tags given by the users. We do some pre-processings similarly to [12]. Here, we give a summary of this pre-processing. We randomly choose 10,000 samples as the training dataset and another 10,000 samples as the testing dataset. The features we used are extracted using the pre-trained neuron network, MatConvNet[37] and there are totally 4096 features. We discretize each feature into 10 discrete numbers based on their quantile value. We extract all the tags and use word2vector[38] technology to map each tag into a low dimensional unit vector. Then, we use the Euclidean distance between those vectors as the distance between tags, and carry out the feature selection on the training dataset. The classifier we use is kNN and the loss function is the top-k cost as the same in the previous experiment. Figure(7) shows the top-5 loss under different numbers of features. From the figure we can see, when the number of features is relatively small, say below 10%, our Wasserstein-distance-based algorithm has less top-5 loss than the KL-distance-based algorithm. Which indicates our algorithm has the power to learn more important features. When the number of features continues increasing, the loss for both of the algorithms increase. It is due to overfitting considering the number of features is large and the training samples is not so many.
Figure 8: The result using the 20 newsgroups dataset.
The third data set we used is the 20 newsgroups dataset[39]. It is a collection of 20,000 documents across 20 different newsgroups. The topics of the 20 newsgroups are summarized in Table(II). In this table, the topics are partitioned according to subject matter. Within each individual cell the topics are very closely related, while between two different cells the topics are unrelated. With the advantage of topic diversity, we set the class distance to be $0.2$ within each topic cell and class distance to be $1$ between different topic cells. The features of the dataset is the top 20000 vocabularies ordered by the term frequency. The value of each feature is computed by uniformly discretizing its tf-idf value into 10 digital numbers. Then we do feature selection with two different algorithms after this pre-processing.
Figure(8) shows the results of two different algorithms. Since the dataset is single labeled, we use the top-1 cost as the loss function and the classifier is kNN as the previous experiments. As we can see, our algorithm provides smaller loss at most cases.
VII Summary
We introduce the Wasserstein distance into one of the traditional feature selection frameworks and demonstrate its potentials in feature selection.
VIII Future Work
The Wasserstein distance based feature has a promising prospect. However, the distance matrix used in the algorithm is hard to define and it varies from datasets to datasets. One future work is to find a unified way to calculate the class distance.